Β Let 𝑓(π‘₯) = ࡝ 1 + 2π‘₯, π‘₯ ≀ 0 3π‘₯ βˆ’ 2,0 < π‘₯ ≀ 1 2π‘₯ ΰ¬Ά βˆ’ 1, π‘₯ > 1 i) Check whether f is discontinuous. If yes, find where? ii) Give a rough sketch of the graph of f.


1
Expert's answer
2021-11-22T07:41:29-0500
"f(x) = \begin{cases}
 1+2x &x\leq 0 \\
 3x-2 &0<x\leq 1\\
2x-1 & x>1
\end{cases}"

i)


"\lim\limits_{x\to 0^-}f(x)=\lim\limits_{x\to 0^-}(1+2x)=1+2(0)=1"

"\lim\limits_{x\to 0^+}f(x)=\lim\limits_{x\to 0^+}(3x-2)=3(0)-2=-2"

"\lim\limits_{x\to 0^-}f(x)=1\not=-2=\lim\limits_{x\to 0^+}f(x)"

"\lim\limits_{x\to 0}f(x) \text{does not exist}"

The function "f(x)" has an jump discontinuity at "x=0."



"\lim\limits_{x\to 1^-}f(x)=\lim\limits_{x\to 1^-}(3x-2)=3(1)-2=1"

"\lim\limits_{x\to 1^+}f(x)=\lim\limits_{x\to 1^+}(2x-1)=2(1)-1=1"

"\lim\limits_{x\to 1^-}f(x)=1=\lim\limits_{x\to 1^+}f(x)=>\lim\limits_{x\to 1}f(x)=1"

"f(1)=3(1)-2=1=\lim\limits_{x\to 1}f(x)"

The function "f(x)" is continuous at "x=1."


The function "f(x)" is discontinuous at "x=0."

The function "f(x)" has an jump discontinuity at "x=0."


ii)